# ME-IGM: Individual-Global-Max in Maximum Entropy Multi-Agent Reinforcement Learning

### v3

@@ Acknowledgment
Table 5: Hyperparameters used for hyper-networks.
Appendix D D. Comparison with Baselines
@@
2. Actor-critic methods utilize the KL divergence as the loss function, while our approach uses the MSE loss. In the CTDE framework, all methods maintain a more accurate centralized Q-function alongside multiple imprecise local policies. The core idea is to distill knowledge from the centralized Q-function to the local policies. Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits. (Kim et al. 2021) pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
@@
4. In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods. However, through experiments, we have demonstrated that our method surpasses both FOP (Zhang et al. 2021) and HASAC (Liu et al. 2023).
@@
mSAC (Pu et al. 2021) introduces a method similar to MASAC. It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
FOP (Zhang et al. 2021) shares a similar issue with (Pu et al. 2021), in that the AC framework it utilizes can actually train the critic with global state information. Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX (Wang et al. 2020). (Hu et al. 2021) experimentally shown that QMIX outperforms QPLEX. Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases. In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
HASAC (Liu et al. 2023), in contrast to FOP (Zhang et al. 2021), avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition. However, our primary contribution addresses the misalignment problem, which is orthogonal to HASAC’s focus. Moreover, since HASAC does not rely on either the IGM or IGO assumptions, establishing a direct theoretical comparison is nontrivial. Nevertheless, we empirically demonstrate that our method consistently outperforms HASAC across benchmark tasks presented in the paper.
PAC (Zhou, Lan, and Aggarwal 2022) utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
Appendix E E. Hyper-parameters
For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication. We used FOP by  Zhang et al. and HASAC by Zhong et al.. Our method, ME-IGM is built on Hu et al.. The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU. The following are the hyperparameters used in the experiments.
@@
[…]
@@
Table 6: HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Liu et al. 2023).
@@
Table 7: ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
@@
Table 8: ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 (Ellis et al. 2022).
@@
Table 9: FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper (Zhang et al. 2021).
@@
Table 10: ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).
@@
Table 11: ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked (Carroll et al. 2019).

### v4

@@ Appendix C Hyper-network Structure
Table 5. Hyperparameters used for hyper-networks.
Appendix D Comparison with Baselines
@@
2. Actor-critic methods utilize the KL divergence as the loss function, while our approach uses the MSE loss. In the CTDE framework, all methods maintain a more accurate centralized Q-function alongside multiple imprecise local policies. The core idea is to distill knowledge from the centralized Q-function to the local policies. Actor-critic methods typically minimize the KL divergence between the actor and the softmax critic, whereas our method minimizes the MSE loss between their logits. kim2021comparing pointed out that, in distillation contexts, MSE is a superior loss function compared to KL divergence.
@@
4. In fact, it’s challenging to theoretically prove that value-based methods are inherently superior to AC methods. However, through experiments, we have demonstrated that our method surpasses both FOP zhang2021fop and HASAC liu2023maximum.
@@
mSAC pu2021decomposed introduces a method similar to MASAC. It employs the AC architecture, thus encountering issues 1 and 2 mentioned above, but does not fully utilize global information to train the critic, hence lacking the advantage noted in point 3.
FOP zhang2021fop shares a similar issue with pu2021decomposed, in that the AC framework it utilizes can actually train the critic with global state information. Furthermore, FOP employs a credit assignment mechanism similar to that of QPLEX wang2020qplex. hu2021rethinking experimentally shown that QMIX outperforms QPLEX. Additionally, FOP’s proof concept is that under the IGO conditions, if we locally move local policies towards local optimal policies, the distance between the joint policy and the optimal joint policy also decreases. In contrast, our proof approach first establishes that our method is equivalent to the single-agent SAC algorithm and then completes the proof using existing conclusions.
HASAC liu2023maximum, in contrast to FOP zhang2021fop, avoids the restrictive Individual-Global-Optimal (IGO) assumption used for joint policy factorization by employing a theoretically grounded sequential optimization of individual policies via joint soft policy decomposition. However, our primary contribution addresses the misalignment problem, which is orthogonal to HASAC’s focus. Moreover, since HASAC does not rely on either the IGM or IGO assumptions, establishing a direct theoretical comparison is nontrivial. Nevertheless, we empirically demonstrate that our method consistently outperforms HASAC across benchmark tasks presented in the paper.
PAC zhou2022pac utilizes global information during centralized training, possessing the advantages noted in point 3. However, it still encounter issues 1 and 2 mentioned above.
Appendix E Hyper-parameters
For all baseline algorithms, we implemented them using the corresponding open-source frameworks and chose the hyperparameters provided by these frameworks for replication. We used FOP by zhang2021fop and HASAC by zhong2023heterogeneousagent. Our method, ME-IGM is built on hu2021rethinking. The experiments were conducted on a computer equipped with 92 GB of RAM, a 40-core CPU, and a GeForce RTX 2080 Ti GPU. The following are the hyperparameters used in the experiments.
@@ Appendix E Hyper-parameters
$\text{auto\_}\alpha$ True
$\alpha\text{\_lr}$ 0.0003
@@
Table 6. HASAC hyperparameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper liu2023maximum.
@@
Table 7. ME-QMIX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 ellis2022smacv2.
@@
Table 8. ME-QPLEX hyperparameters used for SMACv2. We utilize the hyperparameters used in SMACv2 ellis2022smacv2.
@@
Table 9. FOP hyper-parameters used for SMACv2. We use the hyperparameters for SMAC as specified in the original paper zhang2021fop.
@@
Table 10. ME-QMIX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked carroll2019utility.
@@
Table 11. ME-QPLEX hyperparameters used for Overcooked. We utilize the hyperparameters used in Overcooked carroll2019utility.